MONT 100N – Modeling the World Student-Led Discussions on Collapse Week of December 4

General Information

For the final week of the semester, we will continue to think about various aspects of Collapse by Jared Diamond, but different groups in the class will be responsible for coming up with discussion questions or other ways to frame what we do in class on Monday, December 4, Wednesday, December 6 (after the first 20 minutes or so, when the CEFs – course evaluation forms – will be administered), and Friday December 8. The idea is to come up with some questions or approaches to thinking about material from the book in ways that will stimulate participation and interest. Drawing connections with other topics we have discussed will be a plus! Feel free to get creative with visual aids, videos, etc. if you think they will add to the discussion. Everyone in each group should actively participate both in deciding what you want to address and orally in your class meeting. When there are two groups on a day, each is responsible for about 25 minutes of the class.

December 4 – The Case of Norse Greenland

Group 1 – Madison, Eve, Andrew, Logan – Chapter 7 (Norse Greenland’s Flowering)

Some key points to think about:

When did Norse settlers arrive on Greenland? Why did they settle there? What does Diamond say about the climate of Greenland when they arrived? Did the climate stay the same? How did the settlers think of themselves and their relationship with the Native American (Inuit) population of Greenland?

Group 2 – Najee, Emily, Joe, Matthew – Chapter 8 (Norse Greenland’s End)

Some key points to think about:

What is the surviving evidence of how the Greenland settlements came to an end? What factors does Jared Diamond blame for their eventual failure? How did the attitudes of the Norse settlers contribute to this? Could the Norse have succeeded in Greenland if they were willing to change their attitudes? Why were they unable to do that?

December 6

Group 3 – Grace, Genevieve, Lily – Chapter 14 (Why do Some Societies Make Disastrous Decisions?)

Some key points to think about:

Focus on the discussion of the different ways societies have failed to make good decisions about their future. Ask the class whether they can see parallels with things that are going on in the world today.

December 8

Group 4 – Ryan, Connor, Liz – Chapter 15 (Big Businesses and the Environment: Different Conditions, Different Outcomes)

Some key points to think about:

I think it would be interesting to make a contrast here between Diamond’s examples of “good” business behavior and the generally anti-business attitudes that Pope Francis expressed in Laudato Si’. Who’s right? Both? Neither? Is it possible for businesses pursuing profits to be good environmental stewards too?

Group 5 – Christian, Angel, Kyle – Chapter 16 (The World as a Polder)

Some key points to think about:

(What’s a “polder?” Why did Diamond use this for the title of his final chapter?) What are the most serious problems Diamond lists? Why does he reject the “one-liner” objections? It might be interesting to see whether people believe some of them by asking. Why is Diamond moderately hopeful for the future? Do you feel the same way?